# Anisotropic anomalous diffusion in microgravity dusty plasma. Part One: Nonequilibrium Statistical Analysis

### v2

@@ 1 Introduction
Complex (or dusty) plasmas are a collection of electrons, ions, neutral particles, and dust grains (typically micro- to nano-meter in size). Dusty plasma is ubiquitous in astrophysical and space environments, as well as in laboratory settings, both on Earth and in microgravty. Dusty plasma is a unique analogue system for the study of complex phenomena such as phase transitions, anomalous diffusion, and metastability. The particles in these systems are visible at the kinetic level, thus allowing for a reconstruction of the entire phase space. In addition, particle tracking or velocimetry techniques can be used to obtain statistically significant amount of data. Finally, dusty plasma experiments are reasonably simple to build (table-top) and highly controlled, which makes them ideal for deployment in space. Due to the complex interactions among the different charged species, dusty plasmas are observed to exhibit various waves, instabilities, and nonlinear structures [1, 2]. Dusty plasmas are also ideal for studying solid-liquid phase transitions [3, 4, 5, 6], electroheology [7, 8], strong interparticle coupling and long-range interactions [9, 10, 11], kinetic theories and diffusion properties [12, 13, 14, 15, 16], and critical phenomena such as melting and crystallization [6, 16, 17], and turbulence [18, 19, 20]. As they exhibit many-body effects, dusty plasmas are useful analogue systems for the study of complex systems such as condensed matter [21, 18] and smart materials [22, 7, 23]. Many other physical aspects and applications of dusty plasma have been summarized in several recent overview papers[24, 25, 26]. With all the characteristics described above, dusty plasmas are ideal for testing new analytical models, especially those focused on non-equilibrium or non-extensive statistics, anomalous diffusion, and stochasticity [18, 27, 28, 29, 30, 14, 15].
@@
Here we investigate dusty plasma experiments conducted in the Plasmakristall-4 (PK-4) facility on board the International Space Station, where the microgravity environment allows to neglect gravity and confinement forces, thus, focusing on plasma-mediated dust-dust interactions. Recent studies using the PK-4 facility [31] have investigated various dynamical phenomena, including dust ionization waves [32], ion density waves [33], dust acoustic waves [34],unsteady shear flows and flow patterns fluctuations [35], and lunar atmosphere dust [36]. Additionally, analysis of PK-4 data has inspired a breadth of numerical studies, including particle-in-cell simulation of PK-4 predicting the formation of ionization waves [37] and molecular dynamics simulations of dust and ions investigating how such ionization waves can cause anisotropies in the ion wakefields around the dust grains [38, 39].
@@
The PK-4 experiment uses video cameras to track individual dust particles, which allows for obtaining large amounts of statistical information on the dust particles positions and velocities. Previous studies of dusty plasma with the PK-4 experiment have shown velocity distribution functions (VDFs) that were non-Maxwellian, had high-energy tails, and displayed anomalous diffusion. Anomalous diffusion is a microscopic process which leads to a mean squared displacement (MSD) that grows non-linearly with time $MSD\propto\tau^{\alpha}$, where $\tau$ is time delay. If the MSD growth is faster than linear with time, $\alpha>1$, the particles are superdiffusive, while growth rate slower than linear, $\alpha<1$, indicates subdiffusion. Anomalous diffusion and corresponding non-linear MSDs have been observed in biological molecular transport [40], Lévy flight movement patterns of living organisms [41], condensed matter physics [42], and dusty plasmas [43, 14, 30]. A review of classical and anomalous diffusion across many fields can be found in [44]. A major goal of the present paper is to promote the use of Tsallis nonequilibrium statistics in dusty plasma analysis, since it provides greater clarity to the classification and study of nonequilibrium systems, non-Maxwellian distributions, and the resulting anomalous diffusion.
@@
Here we analyze anomalous diffusion and thermodynamic nonequilibrium properties by constructing MSD plots and histograms of dust displacements and velocities. Nonextensive (Tsallis) statistics[45, 46, 47, 48, 49, 50, 51] is used to quantify how dust PDFs deviate from a Gaussian/Maxwellian distribution and to identify the sub-regime of anomalous diffusion that best describe each dataset. Nonextensive statistics (also called nonequilibrium statistics) is a formulation of statistical mechanics where entropies are nonadditive (nonextensive). Nonextensive statistics has been used in plasma physics to better understand solar wind turbulence and dynamics [52, 53], plasma waves [54, 55, 56], collisions [57], dusty plasmas [58, 59], including analysis of previous PK-4 experiments using combined DC-RF discharge [30, 15, 14]. We build on these previous studies by analyzing PK-4 data from pure DC discharge experiments where the dust clouds was shown to exhibit strongly non-isotropic filamentary structure.
@@
Fits to the MSDs plots reveal a non-linear relation with time delay $\tau^{\alpha}$, which is indicative of anomalous diffusion. Fits to both the position and velocity histograms for the direction along the external electric field are best described by a q-Gaussian distribution function instead of a Gaussian or a Maxwellian one. We use scaling relations proposed in literature to compare the nonextensive $q_{p}$ exponent found from fits to displacement histograms against the MSD exponent $\alpha$ to validate the sub-regime of observed anomalous diffusion. The dust displacements and velocities in the direction perpendicular to the external electric field are best described by a Bi-q-Gaussian distribution. This suggests that the system has two distinct thermodynamic populations. As the neutral gas pressure in dusty plasma mediates the dust-neutral collisions, increasing pressure in these experiments acts as decreasing temperature, which is why the observed thermodynamics is sensitive to pressure changes. This was also seen by [61], which conducted a structural analysis of the same set of experiments using pair correlation function techniques. The remainder of this paper is dedicated to the qualitative and quantitative description of the observed anomalous diffusion, anisotropies, and equilibrium properties of dusty plasma in PK-4.
@@ 2 Experimental Setup
The PK-4 Campaign 7 experiments discussed here were conducted on July 26, 2019. Our analysis uses the video data from nine sets of pressure-current conditions (see Table 1). The current polarity switching was 500 Hz with a duty cycle of $50\%$, and microparticle size of $3.38\mu m$ diameter. In each case, the dust cloud was allowed to settle for  50 s, after which a scan of the laser sheet was performed through the dust cloud (along the y-axis in 1). These y-scans allow for obtaining information on the 3D structure of the cloud. The statistical analysis presented here uses particle tracking data obtained for a period in which the dust cloud has settled and that the cameras and laser sheet are focused on the mid-plane of the cloud. We use the dust particle positions and velocities to study the probability distribution functions, anomalous diffusion, temperature, and nonequilibrium properties.
@@ 3.1 Diffusion
where $\Delta=\frac{\partial^{2}}{\partial x^{2}}$ is the Laplacian operator, D is the diffusion constant, and p(x,t) is the distribution function. Equation 1 is also the linear Fokker-Planck equation with no drift. The one-dimensional solution to the diffusion equation has a Gaussiona distribution functional form given by
$p(x,t)=\frac{1}{\sqrt{2\pi Dt}}e^{-\frac{(x-x_{0})^{2}}{2Dt}}.$ (2)
@@
where $d$ is the dimension and $\alpha$ is an exponent that quantifies deviations from a linear dependence in time (i.e., the presence of anomalous diffusion. When $\alpha=1$ (linear MSD), the diffusion is classical, while $\alpha\neq 1$ implies anomalous diffusion. While anomalous diffusion is ubiquitous in nature [40, 41, 42, 43, 14, 30, 44], its mathematical description is challenging. One way to study anomalous diffusion is with a nonlinear diffusion equation, the nonlinear Fokker-Planck differential equation, which has solutions including Porous Medium equations (see page 249 of [62]) and Tsallis q-Gaussian distributions [63] (also known as Kappa distribution or Student’s t distribution). Anomalous diffusion may also be studied by use of fractional spatial derivatives [27, 29, 64, 18] which may have solutions such as Lévy distributions. We note that it is not well understood if universal scaling relations exist that link the different proposed analytical formulations for anomalous diffusion. While there are some direct relations between the nonlinear Fokker-Planck equations and the q-Gaussian solutions, their relation to fractional differential equations is not straightforward even though they can both describe similar regimes of anomalous diffusion.
@@ 3.2 The Nonlinear Diffusion Equation and Tsallis’ Statistics
has the solution
$p(x,t)=\frac{1}{\sqrt{\pi A_{q}}}[1+(q-1)\frac{(x-x_{0})^{2}}{A_{q}(D_{q}t)^{\frac{1}{3-q}}}]^{\frac{-1}{q-1}}.$ (5)
Here $\nu$ is related to the nonextensive parameter $q$ by $\nu\equiv 2-q$ [63]. $A_{q}$ is a normalization factor given by
$A_{q}=\begin{cases}\sqrt{(q-1)}\frac{\Gamma(\frac{1}{q-1})}{\Gamma(\frac{3-q}{2q-2})}&\text{if }1<q<3\\ 2&\text{if }q=1\\ \sqrt{(1-q)}\frac{\Gamma(\frac{5-3q}{2-2q})}{\Gamma(\frac{2-q}{1-q})}&\text{if}q<1.\end{cases}$ (6)
@@
The parameter $q$ can also be related to the distribution kurtosis $\kappa$, through the following relationship: $k=\frac{15-9q}{7-5q}$, for $q<7/5$ [65] where $k=Kurtosis=\frac{\langle x^{4}\rangle}{\langle x^{2}\rangle^{2}}$. This can be used to quantify the deviations from a Gaussian distribution and the related anomalous diffusion (e.g., as discussed in [66]). The primary interest of the present study is the range $31<q<$ 3 where the distributions exhibit leptokurtic or ‘fat-tail’ behavior. This regime is also where anomalous diffusion is superdiffusive ($\alpha>1$), though we will also discuss subdiffusive cases where $q<1$ or $\alpha<1$ (see Figure 2 for example distributions representing each regime). Anomalous diffusion is described by taking the second moment of Eq. 5, which yields a non-linear relation between the MSD and time
$\langle(x-x_{0})^{2}\rangle\propto D_{q}t^{\frac{2}{3-q}};\hskip 17.07164ptq<\frac{5}{3},$ (7)
@@
where $D_{q}$ is a new diffusion coefficient. Nonextensive statistics were proposed by Tsallis [67] and are described fully in the textbook on the subject [63]. The probability distribution function in this formulation is called a q-Gaussian, where the parameter $q$ quantifies the nonextensivity of the system. The nonextensivity feature of the formulation makes it independent of initial conditions and suitable for modeling many-body complex systems with long-range interactions, such as plasma. Systems with this nonextensive behavior no longer have additive entropies, $S_{q}(A+B)\neq S_{q}(A)+S_{q}(B)$, and the different microstates are not in equilibrium. Therefore, quantifying the nonextensivity is needed to make rigorous statistical claims and understand equilibrium properties of systems with long-range interactions and correlations. Thus, $q$ can be understood as a parameter quantifying the strength of correlations or non-local interactions causing the system to move away from equilibrium. The nonextensive parameter $q$ can be found by extremizing the Tsallis entropy $S_{q}$ with Lagrange constraints, such as normalization and momement of the distribution. Essentially, it can be thought of as the nonlinear form of the Maxwellian distribution, with normalization $A_{q}$, kinetic energies $E_{i}$, potential $U$, and inverse temperature $\beta_{q}$. This results in the following probability distribution function, known as the q-Gaussian
@@
The relationship between the Maxwellian temperature $T_{M}$ and the q-Gaussian temperature $T_{q}$ was discussed in [56] and is given by
@@ 4.1 Particle Tracking and Drift Subtraction
Figure 3 shows the case where the most pronounced nonhomogeneous drift was observed for 70 Pa pressure and 0.35 mA dc current. As can be seen in Fig. 3 B, in the presence of nonhomogeneous drift, applying the drift subtraction to the large region of interest produces particle trajectories that exhibit some coordinated motion (instead of showing random diffusive motion). The NHDS method seems to work better as shown in 3 C, though some small areas still exhibit coordinated motion instead of the desired diffusive behavior. The drifts for each smaller domain region are shown in 3 D. While the total drift simply implied that the dust particles were linearly drifting to the left, the domain drifts show a spread of drift directions.
@@
As a quantitative check of the NHDS application of drift correction for the 70 Pa 0.35 mA case, we calculated the velocity autocorrelation of all domains before and after, shown in Fig. 4. While the velocity autocorrelation before subtraction (red) showed a nonzero value, after the NHDS subtraction (blue), the total autocorrelation dropped to zero. Using these drift-subtracted particle trajectories, we proceeded to find the values of $\alpha$ from a fit to the MHDS plots and the values of $q_{p}$ and $q_{v}$ from fits to the position and velocity histograms. The values of $\alpha$ obtained from fits to the MSD plots at time delay of five seconds are very similar for two drift subtraction methods. However, the application of the NHDS method decreased the magnitude of th measured MSD (reflected in the vertical axis) by an order of magnitude.

### v3

# Anisotropic anomalous diffusion in microgravity dusty plasma. Part One: Nonextensive Statistical Analysis

@@ 1 Introduction
Complex (or dusty) plasmas are a collection of electrons, ions, neutral particles, and dust grains (typically micro- to nano-meter in size). Dusty plasma is ubiquitous in astrophysical and space environments, as well as in laboratory settings, both on Earth and in microgravty. Dusty plasma is a unique analogue system for the study of complex phenomena such as phase transitions, anomalous diffusion, and metastability. The particles in these systems are visible at the kinetic level, thus allowing for a reconstruction of the entire phase space. In addition, particle tracking or velocimetry techniques can be used to obtain statistically significant amount of data. Finally, dusty plasma experiments are reasonably simple to build (table-top) and highly controlled, which makes them ideal for deployment in space. Due to the complex interactions among the different charged species, dusty plasmas are observed to exhibit various waves, instabilities, and nonlinear structures [1, 2].
Dusty plasmas are also ideal for studying solid-liquid phase transitions [3, 4, 5, 6], electroheology [7, 75, 76], strong interparticle coupling and long-range interactions [8, 9, 10], kinetic theories and diffusion properties [11, 12, 13, 14], and critical phenomena such as melting and crystallization [6, 14, 15], and turbulence [16, 17, 18]. As they exhibit many-body effects, dusty plasmas are useful analogue systems for the study of complex systems such as condensed matter [19, 16] and smart materials [20, 7, 21]. Many other physical aspects and applications of dusty plasma have been summarized in several recent overview papers[22, 23, 24]. With all the characteristics described above, dusty plasmas are ideal for testing new analytical models, especially nonequilibrium statistical mechanics anomalous diffusion, and stochasticity [16, 25, 26, 27, 28, 13]. In part two of this paper, we will be analyzing a very recent spectral model using an Anderson Type Hamiltonian with a long-range Fractional Laplacian operator [16, 25, 26, 27] and relating it to the same data set presented here.
@@
Here we investigate dusty plasma experiments conducted in the Plasmakristall-4 (PK-4) facility on board the International Space Station, where the microgravity environment allows to neglect gravity and confinement forces, thus, focusing on plasma-mediated dust-dust interactions. Recent studies using the PK-4 facility [29] have investigated various dynamical phenomena, including dust ionization waves [30, 77, 78], ion density waves [31], dust acoustic waves [32],unsteady shear flows and flow patterns fluctuations [33], and lunar atmosphere dust [34]. Additionally, analysis of PK-4 data has inspired a breadth of numerical studies, including particle-in-cell simulation of PK-4 predicting the formation of ionization waves [35] and molecular dynamics simulations of dust and ions investigating how such ionization waves can cause anisotropies in the ion wakefields around the dust grains [36, 37]. Finally, non-Mazwellian distributions of the dust velocities have been observed in PK-4 experiments with RF discharge configuration [13]. Here we discuss the observation of non-Maxwellian distribution functions in PK-4 experiments with a pure DC discharge, where a polarity-switching external electric field causes anisotropies on the dust-dust interaction potential and the resulting dust diffusion.
@@
The PK-4 experiment uses video cameras to track individual dust particles, which allows for obtaining large amounts of statistical information on the dust particles positions and velocities. Previous studies of dusty plasma with the PK-4 experiment have shown velocity distribution functions (VDFs) that were non-Maxwellian, had high-energy tails, and displayed anomalous diffusion. Anomalous diffusion is a microscopic process which leads to a mean squared displacement (MSD) that grows non-linearly with time $MSD\propto\tau^{\alpha}$, where $\tau$ is time delay. If the MSD growth is faster than linear with time, $\alpha>1$, the particles are superdiffusive, while growth rate slower than linear, $\alpha<1$, indicates subdiffusion. Anomalous diffusion and corresponding non-linear MSDs have been observed in biological molecular transport [38], Lévy flight movement patterns of living organisms [39], condensed matter physics [40], and dusty plasmas [41, 28]. A review of classical and anomalous diffusion across many fields can be found in [42]. One of the goals of the present paper is to promote the use of Tsallis nonextensive statistics in dusty plasma analysis, since it provides greater clarity to the classification and study of nonequilibrium systems, non-Maxwellian distributions, and the resulting anomalous diffusion.
@@
Here we analyze anomalous diffusion and thermodynamic nonequilibrium properties by constructing MSD plots and histograms of dust displacements and velocities. Nonextensive (Tsallis) statistics [43, 44, 45, 46, 47, 48, 49] is used to quantify how dust PDFs deviate from a Gaussian/Maxwellian distribution and to identify the sub-regime of anomalous diffusion that best describe each dataset. Nonextensive statistics is a formulation of statistical mechanics where entropies are nonadditive (nonextensive). Nonextensive statistics has been used in plasma physics to better understand solar wind turbulence and dynamics [50, 51], plasma waves [52, 53, 54], collisions [55], dusty plasmas [56, 57], including analysis of previous PK-4 experiments using combined DC-RF discharge [28, 13]. We build on these previous studies by analyzing PK-4 data from pure DC discharge experiments where the dust clouds were shown to exhibit strongly non-isotropic filamentary structure.
@@
Fits to the MSDs plots reveal a non-linear relation with time delay $\tau^{\alpha}$, which is indicative of anomalous diffusion. Fits to both the position and velocity histograms for the direction along the external electric field are best described by a q-Gaussian distribution function, where the non-extensive exponent $q$ quantifies the ’tailedness’ of a distribution deviates from a standard Gaussian or a Maxwellian one. To determine the diffusion sub-regime, we use scaling relations from literature to compare the nonextensive exponent $q_{p}$ of position distributions (found from fits to displacement histograms) against the MSD exponent $\alpha$. The dust displacements and velocities in the direction perpendicular to the external electric field are best described by a Bi-q-Gaussian distribution, which is a sum of two q-Gaussian distributions. This suggests that the system has two distinct thermodynamic populations. As the neutral gas pressure in dusty plasma mediates the dust-neutral collisions, increasing pressure in these experiments acts as decreasing temperature, which is why the observed thermodynamics is sensitive to pressure changes. This was also seen by [59], which conducted a structural analysis of the same set of experiments using pair correlation function techniques. The remainder of this paper is dedicated to the qualitative and quantitative description of the observed anomalous diffusion, anisotropies, and equilibrium properties of dusty plasma in PK-4.
@@ 2 Experimental Setup
The PK-4 Campaign 7 experiments discussed here were conducted on July 26, 2019. Our analysis uses the video data from nine sets of pressure-current conditions (see Table 1). The current polarity switching was 500 Hz with a duty cycle of $50\%$, and microparticle size of $3.38\mu m$ diameter. In each case, the dust cloud was allowed to settle for  50 s, after which a scan of the laser sheet was performed through the dust cloud (along the y-axis in Figure 1). These y-scans allow for obtaining information on the 3D structure of the cloud. The statistical analysis presented here uses particle tracking data obtained for a period in which the dust cloud has settled and that the cameras and laser sheet are focused on the mid-plane of the cloud. We use the dust particle positions and velocities to study the probability distribution functions, anomalous diffusion, temperature, and nonequilibrium properties.
@@ 3.1 Diffusion
where $\Delta=\frac{\partial^{2}}{\partial x^{2}}$ is the Laplacian operator, D is the diffusion constant, and p(x,t) is the distribution function. Equation (1) is also the linear Fokker-Planck equation with no drift. The one-dimensional solution to the diffusion equation has a Gaussion distribution functional form given by
$p(x,t)=\frac{1}{\sqrt{4\pi Dt}}e^{-\frac{(x-x_{0})^{2}}{4Dt}}.$ (2)
@@
where $d$ is the dimension and $\alpha$ is an exponent that quantifies deviations from a linear dependence in time (i.e., the presence of anomalous diffusion. When $\alpha=1$ (linear MSD), the diffusion is classical, while $\alpha\neq 1$ implies anomalous diffusion. While anomalous diffusion is ubiquitous in nature [38, 39, 40, 41, 28, 42], its mathematical description is challenging. One way to study anomalous diffusion is with a nonlinear diffusion equation, the nonlinear Fokker-Planck differential equation, which has solutions including Porous Medium equations (see page 249 of [60]) and Tsallis q-Gaussian distributions [61] (also known as Kappa distribution or Student’s t distribution). Anomalous diffusion may also be studied by use of fractional spatial derivatives [25, 27, 62, 16] which may have solutions such as Lévy distributions. We note that it is not well understood if universal scaling relations exist that link the different proposed analytical formulations for anomalous diffusion. While there are some direct relations between the nonlinear Fokker-Planck equations and the q-Gaussian solutions, their relation to fractional differential equations is not straightforward even though they can both describe similar regimes of anomalous diffusion.
@@ 3.2 The Nonlinear Diffusion Equation and Tsallis’ Statistics
has the following explicit solution derived in [60], page 249
$p(x,t)=A_{\nu}^{-1}\left[2D(\nu+\nu^{2})(t-t_{0})\right]^{\frac{-1}{1+\nu}}\cdot\left[1+\frac{(1-\nu)(x-x_{0})^{2}}{\left((2D(\nu+\nu^{2})(t-t_{0})A_{\nu}^{1-\nu})^{\frac{2}{1+\nu}}\right)}\right]^{\frac{-1}{1-\nu}}.$ (5)
The expression can be rewritten using Tsallis nonextensive $q$ using the substitution $\nu\equiv 2-q$, which yields
$p(x,t)=A_{q}^{-1}\left[2D(6-5q+q^{2})(t-t_{0})\right]^{\frac{-1}{3-q}}\left[1+\frac{(q-1)(x-x_{0})^{2}}{\left((2D(6-5q+q^{2})(t-t_{0})A_{q}^{q-1})^{\frac{2}{3-q}}\right)}\right]^{\frac{-1}{q-1}}$ (6)
This is similar to the expression provided in [61], where $A_{q}$ is a normalization factor given by
$A_{q}=\begin{cases}\sqrt{\frac{\pi}{(1-q)}}\frac{\Gamma(\frac{2-q}{1-q})}{\Gamma(\frac{q}{2-2q})}&\text{if }q<1\\ \sqrt{\pi}&\text{if }q=1\\ \sqrt{\frac{\pi}{(q-1)}}\frac{\Gamma(\frac{3-q}{2q-2})}{\Gamma(\frac{q}{q-1})}&\text{if }1<q<3.\end{cases}$ (7)
@@
When $q=1$ and $t_{0}=0$ this recovers equation 2. The parameter $q$ can also be related to the distribution kurtosis $k$, through the following relationship: $k=\frac{15-9q}{7-5q}$, for $q<7/5$ [63] where $k=Kurtosis=\frac{\langle x^{4}\rangle}{\langle x^{2}\rangle^{2}}$. This can be used to quantify the deviations from a Gaussian distribution and the related anomalous diffusion (e.g., as discussed in [64]). The primary interest of the present study is the range $31<q<$ 3 where the distributions exhibit leptokurtic or ‘fat-tail’ behavior. This regime corresponds to superdiffusive transport ($\alpha>1$), though we will also discuss subdiffusive cases where $q<1$ or $\alpha<1$ (see Figure 2 for example distributions representing each regime). Anomalous diffusion is described by taking the second moment of Eq. 5, which yields a non-linear relation between the MSD and time
$\langle(x-x_{0})^{2}\rangle=B_{q}(D\tau)^{\frac{2}{3-q}},\quad\tau=t-t_{0}$ (8)
where
$B_{q}=\frac{2^{\frac{1-q}{q-3}}\pi^{\frac{3(1-q)}{2(q-3)}}(q-1)^{\frac{q}{2(q-3)}}\Gamma\left(\frac{4-3q}{2q-2}\right)}{(q-3)^{\frac{2}{q-3}}(q-2)^{\frac{2}{q-3}}\Gamma\left(\frac{1}{q-1}\right)\left(\frac{\Gamma\left(\frac{2-q}{2q-2}\right)}{\Gamma\left(\frac{q}{q-1}\right)}\right)^{\frac{4q-6}{q-3}}}$ (9)
@@
Nonextensive statistics was proposed by Tsallis [65] and is fully described in his textbook [61]. The probability distribution function in this formulation is called a q-Gaussian, where the parameter $q$ quantifies the nonextensivity of the system. The nonextensivity feature of the formulation makes it independent of initial conditions and suitable for modeling many-body complex systems with long-range interactions, such as plasma. Systems with this nonextensive behavior no longer have additive entropies, $S_{q}(A+B)\neq S_{q}(A)+S_{q}(B)$, and the different microstates are not in equilibrium. Therefore, quantifying the nonextensivity is needed to make rigorous statistical claims and understand equilibrium properties of systems with long-range interactions and correlations. Thus, $q$ can be understood as a parameter quantifying the strength of correlations or non-local interactions causing the system to move away from equilibrium. The nonextensive parameter $q$ can be found by extremizing the Tsallis entropy $S_{q}$ with Lagrange constraints, such as normalization and moment of the distribution. Essentially, it can be thought of as the nonlinear form of the Maxwellian distribution, with normalization $A_{q}$, kinetic energies $E_{i}$, potential $U$, and inverse temperature $\beta_{q}$. This results in the following probability distribution function, known as the q-Gaussian
@@
The relationship between the Maxwellian kinetic temperature $T_{M}$ and the q-Gaussian kinetic temperature $T_{q}$ was discussed in [54] and is given by
@@ 4.1 Particle Tracking and Drift Subtraction
Figure 3 shows the case where the most pronounced nonhomogeneous drift was observed for 70 Pa pressure and 0.35 mA dc current. As can be seen in Figure 3 B, in the presence of nonhomogeneous drift, applying the drift subtraction to the large region of interest produces particle trajectories that exhibit some coordinated motion (instead of showing random diffusive motion). The NHDS method seems to work better as shown in 3 C, though some small areas still exhibit coordinated motion instead of the desired diffusive behavior. The drifts for each smaller domain region are shown in 3 D. While the total drift simply implied that the dust particles were linearly drifting to the left, the domain drifts show a spread of drift directions.
@@
As a quantitative check of the NHDS application of drift correction for the 70 Pa 0.35 mA case, we calculated the velocity autocorrelation of all domains before and after, shown in Figure 4. While the velocity autocorrelation before subtraction (red) showed a nonzero value, after the NHDS subtraction (blue), the total autocorrelation dropped to zero. Using these drift-subtracted particle trajectories, we proceeded to find the values of $\alpha$ from a fit to the MHDS plots and the values of $q_{p}$ and $q_{v}$ from fits to the position and velocity histograms. The values of $\alpha$ obtained from fits to the MSD plots at time delay of five seconds are very similar for two drift subtraction methods. However, the application of the NHDS method decreased the magnitude of th measured MSD (reflected in the vertical axis) by an order of magnitude.